# AutoMMLab: Automatically Generating Deployable Models from Language Instructions for Computer Vision Tasks

### v1

@@ 7 Broad impact and ethics statement
#Iter
Image classification
Object detection
Semantic segmentation
Keypoint detection
1
0.6704$\pm$0.0609
0.3087$\pm$0.0357
0.6144$\pm$0.0552
0.4570$\pm$0.0667
2
0.8156$\pm$0.0488
0.3866$\pm$0.0210
0.7460$\pm$0.0354
0.6103$\pm$0.0483
3
0.8869$\pm$0.0346
0.4133$\pm$0.0132
0.7960$\pm$0.0245
0.6683$\pm$0.0305
4
0.9200$\pm$0.0270
0.4249$\pm$0.0085
0.8196$\pm$0.0178
0.6932$\pm$0.0203
5
0.9418$\pm$0.0200
0.4310$\pm$0.0059
0.8326$\pm$0.0130
0.7060$\pm$0.0146
6
0.9539$\pm$0.0151
0.4345$\pm$0.0042
0.8398$\pm$0.0088
0.7135$\pm$0.0111
7
0.9614$\pm$0.0120
0.4368$\pm$0.0032
0.8440$\pm$0.0063
0.7183$\pm$0.0090
8
0.9665$\pm$0.0093
0.4384$\pm$0.0022
0.8466$\pm$0.0044
0.7226$\pm$0.0072
9
0.9702$\pm$0.0072
0.4395$\pm$0.0018
0.8482$\pm$0.0033
0.7252$\pm$0.0061
10
0.9723$\pm$0.0064
0.4404$\pm$0.0015
0.8493$\pm$0.0023
0.7277$\pm$0.0055
Appendix A6 Discussions
A6.1 Comparisons with other AutoML systems
@@
AutoMMLab is a research project, whose goal is not to create a commercial system, but to showcase how LLM-empowered AutoML system can achieve fully end-to-end prompt-based model training, which is expected to have broad impact for the community. However, in Table A11, we also compare with several AutoML systems for reference. As far as we know, we are the first to build a request-to-model AutoML system for computer vision. Previous AutoML systems are unable to perform our newly proposed “end-to-end prompt-based model training” task. The right table compares AutoML systems in the following dimensions: Request Understanding (RU), Programming-Free (PF), Data Preparation (DP), Model Preparation (MP), Hyperparameter Optimizer (HPO), Supporting Computer Vision tasks (SCV). ✓ means automatic,  means requiring manual operation, ✗ means not supported. Compared to other systems, AutoMMLab is fully automatic, end-to-end, and easier-to-use.
Table A11: Comparisons with other AutoML systems.
@@
AutoKeras
Amazon SageMaker
Google AutoML
@@
A6.2 Limitations and failure cases
The failure cases mainly result from the misunderstanding of user requirements and the absence of the model or data that meets the requirements. The former can be addressed through an improved RU model, whereas the latter can be remedied by expanding the model zoo and data zoo. Although AutoMMLab looks somewhat rudimentary, we believe it holds significant potential.

### v2

@@ 7 Conclusion
#Iter Image classification Object detection Semantic segmentation Keypoint detection
1 0.6704$\pm$0.0609 0.3087$\pm$0.0357 0.6144$\pm$0.0552 0.4570$\pm$0.0667
2 0.8156$\pm$0.0488 0.3866$\pm$0.0210 0.7460$\pm$0.0354 0.6103$\pm$0.0483
3 0.8869$\pm$0.0346 0.4133$\pm$0.0132 0.7960$\pm$0.0245 0.6683$\pm$0.0305
4 0.9200$\pm$0.0270 0.4249$\pm$0.0085 0.8196$\pm$0.0178 0.6932$\pm$0.0203
5 0.9418$\pm$0.0200 0.4310$\pm$0.0059 0.8326$\pm$0.0130 0.7060$\pm$0.0146
6 0.9539$\pm$0.0151 0.4345$\pm$0.0042 0.8398$\pm$0.0088 0.7135$\pm$0.0111
7 0.9614$\pm$0.0120 0.4368$\pm$0.0032 0.8440$\pm$0.0063 0.7183$\pm$0.0090
8 0.9665$\pm$0.0093 0.4384$\pm$0.0022 0.8466$\pm$0.0044 0.7226$\pm$0.0072
9 0.9702$\pm$0.0072 0.4395$\pm$0.0018 0.8482$\pm$0.0033 0.7252$\pm$0.0061
10 0.9723$\pm$0.0064 0.4404$\pm$0.0015 0.8493$\pm$0.0023 0.7277$\pm$0.0055
A5.1 Experimental Compute Resources and Hyperparameter Settings
The fine-tuning of RU-LlaMA and HPO-LlaMA is executed on 8 NVIDIA Tesla A100 GPUs, each with 80G of GPU memory. The test of AutoMMLab is conducted on 8 NVIDIA Tesla V100 GPUs, each with 32G of GPU memory. When fine-tuning LlaMA2-7b to obtain RU-LlaMA and HPO-LlaMA, we used the AdamW optimizer to train 3 epochs. The initial learning rate was set to 1e-4, and decayed by 0.2 times after each epoch. The maximum text length is set to 4096 and the batch size is set to 4. The rank in LoRA was set to 8 and alpha to 32, fine-tuning the ’q_proj’ and ’v_proj’ layers. The mixed-precision approach is used for acceleration during training. When training multiple rounds of dialogue, only the output of the last round of dialogue is used as supervision information.
A5.2 HPO Experimental Results of Random Sampling
Table A10 presents the detailed quantitative results of hyperparameter optimization with the random sampling baseline. The evaluation metrics employed for image classification, object detection, semantic segmentation, and keypoint detection are accuracy, mean Average Precision (mAP), mean Intersection over Union (mIoU), and mAP, respectively.
A5.3 Detailed HPO Experimental Results on LAMP
Table A13, Table A13, Table A13 and Table A14 respectively list the experimental results of different models on the LAMP hyperparameter optimization dataset. It contains four computer vision tasks(image classification, semantic segmentation, object detection, keypoint detection), each task contains 20 requests equipped with corresponding train dataset, test dataset and model. The table shows the experimental results on the test dataset of models trained using hyperparameter configurations generated by different methods.
Table A11: Detailed results of 20 classification requests. Evaluated by the accuracy on test dataset.
Model 1 2 3 4 5 6 7 8 9 10
BayesianRF 0.333 0.760 0.750 0.960 0.795 0.920 0.760 0.333 0.500 0.640
BayesianGP 0.480 0.987 0.750 0.990 0.860 0.920 0.985 0.553 0.500 0.947
LLaMA2-7B-Chat 0.965 0.333 0.947 0.750 0.960 0.985 0.993 0.990 0.653 0.500
PaLM2 0.975 0.993 0.993 0.750 0.990 1.00 1.00 0.995 0.993 0.950
GPT-3.5-turbo 0.955 0.720 0.973 0.750 0.980 0.970 0.980 0.985 0.333 0.500
GPT-4 0.965 0.333 0.967 0.750 0.960 0.980 0.973 0.975 0.660 0.500
HPO-LLaMA 0.975 0.993 0.973 0.995 0.980 0.995 1.00 0.995 1.00 0.940
Model 11 12 13 14 15 16 17 18 19 20
BayesianRF 0.713 0.333 0.665 0.360 0.490 0.935 0.640 0.675 0.805 0
BayesianGP 0.753 0.333 0.940 0.870 0.945 0.935 0.940 0.710 0.825 0
LLaMA2-7B-Chat 0.907 0.933 0.333 0.915 0.910 0.925 0.960 0.885 0.935 0.990
PaLM2 0.940 0.953 0.993 0.920 0.975 0.945 0.995 0.935 0.990 0.995
GPT-3.5-turbo 0.887 0.987 0.333 0.930 0.960 0.915 0.980 0.880 0.965 1.00
GPT-4 0.900 0.967 0.620 0.915 0.960 0.920 0.975 0.920 0.985 0.995
HPO-LLaMA 0.947 0.967 1.00 0.905 0.985 0.935 1.00 0.935 0.980 1.00
Table A12: Detailed results of 20 object detection requests. Evaluated by the mAP on test dataset.
Model 1 2 3 4 5 6 7 8 9 10
BayesianRF 0.396 0 0.458 0.378 0.442 0 0.002 0.664 0.266 0.217
BayesianGP 0.179 0 0.466 0.377 0.217 0 0.003 0.698 0.269 0.240
LLaMA2-7B-Chat 0 0 0.464 0 0 0.421 0.003 0 0.293 0.233
PaLM2 0.434 0.573 0.472 0.413 0.441 0.451 0.003 0.725 0.304 0.241
GPT-3.5-turbo 0.420 0.568 0.477 0.404 0.456 0.432 0.003 0.712 0.299 0.238
GPT-4 0.413 0.569 0.475 0.414 0.453 0.444 0.140 0.716 0.301 0.237
HPO-LLaMA 0.402 0.569 0.477 0.405 0.445 0.414 0.135 0.724 0.303 0.243
Model 11 12 13 14 15 16 17 18 19 20
BayesianRF 0.483 0.519 0.206 0 0.342 0.558 0.029 0.456 0.275 0.123
BayesianGP 0.499 0.535 0.228 0.094 0.333 0.554 0.033 0.472 0.276 0.125
LLaMA2-7B-Chat 0 0 0.220 0.100 0.371 0 0.030 0 0.301 0.126
PaLM2 0.510 0.540 0.225 0.104 0.378 0.590 0.031 0.477 0.305 0.128
GPT-3.5-turbo 0.511 0.539 0.224 0.104 0.376 0.586 0.031 0.468 0.302 0.128
GPT-4 0.507 0.535 0.217 0.359 0.373 0.590 0.636 0.475 0.296 0.537
HPO-LLaMA 0.507 0.532 0.230 0.372 0.377 0.586 0.660 0.485 0.301 0.539
Table A13: Detailed results of 20 segmentation requests. Evaluated by the mIoU on test dataset.
Model 1 2 3 4 5 6 7 8 9 10
BayesianRF 0.871 0.912 0.815 0.828 0.915 0.802 0.889 0.871 0.869 0.889
BayesianGP 0.861 0.923 0.832 0.828 0.902 0.802 0.889 0.866 0.864 0.891
LLaMA2-7B-Chat 0 0 0 0 0 0 0 0 0.861 0
PaLM2 0.874 0.926 0.934 0.835 0.919 0.809 0.889 0.871 0.864 0.911
GPT-3.5-turbo 0.870 0.927 0.833 0.831 0.920 0.809 0.890 0.875 0.862 0.904
GPT-4-turbo 0 0.921 0.825 0.832 0.905 0.803 0.889 0.874 0.860 0.901
HPO-LLaMA 0.872 0.924 0.830 0.832 0.911 0.813 0.889 0.873 0.868 0.897
Model 11 12 13 14 15 16 17 18 19 20
BayesianRF 0.874 0.765 0.807 0.861 0.828 0.895 0.792 0.820 0.825 0.809
BayesianGP 0.870 0.766 0.811 0.859 0.832 0.897 0.800 0.840 0.827 0.800
LLaMA2-7B-Chat 0.868 0.748 0.763 0.864 0 0.900 0 0 0.822 0
PaLM2 0.872 0.757 0.627 0.864 0.830 0.906 0.796 0.851 0.824 0.842
GPT-3.5-turbo 0.873 0.761 0.797 0.862 0.827 0.897 0.800 0.847 0.827 0.833
GPT-4-turbo 0.870 0.746 0.793 0.861 0.825 0.899 0.799 0.821 0.825 0.814
HPO-LLaMA 0.871 0.763 0.809 0.867 0.828 0.913 0.801 0.849 0.828 0.839
Table A14: Detailed results of 20 keypoint detection requests. Evaluated by the mAP on test dataset.
Model 1 2 3 4 5 6 7 8 9 10
BayesianRF 0.457 0 0.208 0 0.092 0 0 0 0 0
BayesianGP 0 0 0.078 0.173 0 0.118 0.359 0 0 0
LLaMA2-7B-Chat 0 0 0 0 0 0 0 0 0 0
PaLM2 0.883 0.716 0.705 0.686 0.780 0.566 0.733 0.706 0.781 0.698
GPT-3.5-turbo 0.515 0.285 0.192 0.131 0.164 0.392 0.329 0.147 0.148 0.170
GPT-4-turbo 0.218 0.002 0.002 0.000 0.058 0.566 0.003 0.003 0.039 0.262
HPO-LLaMA 0.863 0.741 0.711 0.642 0.743 0.755 0.717 0.686 0.768 0.691
Model 11 12 13 14 15 16 17 18 19 20
BayesianRF 0 0.401 0 0.135 0 0 0 0.004 0.065 0.020
BayesianGP 0 0.249 0 0 0 0.044 0 0.558 0 0.039
LLaMA2-7B-Chat 0 0 0 0 0 0 0 0 0 0
PaLM2 0.788 0.529 0.660 0.758 0.782 0.734 0.790 0.674 0.690 0.726
GPT-3.5-turbo 0.398 0.529 0.102 0.305 0.036 0.028 0.046 0.077 0.078 0.004
GPT-4-turbo 0.016 0.172 0.074 0.419 0.000 0.002 0.000 0.042 0.040 0.000
HPO-LLaMA 0.764 0.786 0.663 0.698 0.764 0.719 0.776 0.677 0.703 0.688
Appendix A6 Limitations
AutoMMLab has presented a general-purpose LLM-empowered AutoML system that follows user’s language instructions to automate the whole model production workflow for computer vision tasks. However there still remain some limitations and improvement spaces: 1) Request Understanding: Whether the generated models can meet the user’s requirements perfectly depends largely on the accuracy of understanding the user’s requests. Understanding the multimodal information input by the users and interacting with the users in a multimodal form can more accurately understand the user’s requests. 2) Data Selection: Our method builds a dataset zoo to find images related to user’s requests and build training datasets. Model generation will fail if no images related to the user’s requests are found in the dataset zoo. Building a more powerful database and generating training data by generating models can effectively improve the success rate of model generation. 3) Model Selection: The models supported by our method are limited to those already available in OpenMMLab. Personalized customized models according to user descriptions can meet more diverse model requirements of users. 4) LLM Hallucination: Although LLMs have the strong generative ability, they possibly still failure to follow instructions present hallucination. How to mitigate hallucination in LLMs and reduce the effects of hallucination should be considered when designing the system.
Appendix A7 Broad impact and ethics statement
The accessibility of any powerful technology to the general public raises ethical considerations (Widder et al. 2022), particularly concerning the potential misuse by malicious actors. The deployment of AutoMMLab poses a risk of generating models that may produce toxic or offensive content when exposed to harmful input. To address these ethical concern, our design emphasizes transparency throughout the entire process and incorporates rigorous security measures in the construction of dataset zoo and model zoo, ensuring the reliability of the generated models.
Appendix A8 Comparisons with other AutoML systems
@@
AutoMMLab is a research project, whose goal is not to create a commercial system, but to showcase how LLM-empowered AutoML system can achieve fully end-to-end prompt-based model training, which is expected to have broad impact for the community. In Table A3, we compare with several AutoML systems for reference. As far as we know, we are the first to build a request-to-model AutoML system for computer vision tasks. Previous AutoML systems are unable to perform our newly proposed “end-to-end prompt-based model training” task.
Figure A3: Comparisons with other AutoML systems in the following dimensions: Request Understanding (RU), Programming-Free (PF), Data Preparation (DP), Model Preparation (MP), Hyperparameter Optimizer (HPO), Supporting Computer Vision tasks (SCV). ✓ means automatic,  means requiring manual operation, ✗ means unsupported.
@@
AutoKeras (Jin et al. 2023)
Amazon SageMaker (Das et al. 2020)
Google AutoML (Bisong and Bisong 2019)
Prompt2Model (Viswanathan et al. 2023b)
@@
Appendix A9 Example of End-to-end Model Production
Figure A4: Example of end-to-end model production workflow.